If Z1 and Z2 are the two complex numbers such that | z1-z2| = | | z1|-|z2| | , then find the principal argument of Z1/ Z2

If Z1 and  Z2 are the two complex numbers such that | z1-z2| = | | z1|-|z2| | , then find the principal argument of Z1/ Z2
 

Grade:11

1 Answers

xxx
5 years ago
If | z1-z2| = | | z1|-|z2| , then the argument of z1= argument of z2( which you can verify using geometry), so arg( z1/z2) = argz1- arg z2= 0